2x+3=5 what is value of x
1
Expert's answer
2020-01-10T10:16:55-0500

"2x+3=5"

Subtract 3 from the left and right sides of the equation

"2x+3-3=5-3"

"2x=2"

Divide the left and right sides by 2,

"x=1"